OpenStudy (anonymous):

John made a rectangle pen for his dog using 28 feet of fencing. If the width of the pen is 2 feet more than one-half the length, what is the length and width of the pen? The length of the pen is

OpenStudy (anonymous):

let L = length and W = width. Perimeter = 28 = 2L + 2W and W = L + 2 plug that into the first equation: 2L + 2(L + 2) = 28 solve for L

OpenStudy (anonymous):

is it 24

OpenStudy (anonymous):

4L = 24 divide both by 4

OpenStudy (anonymous):

i ya thats what i forgot lol

OpenStudy (anonymous):

6